# Can AI Prompt Humans? Multimodal Agents Prompt Players’ Game Actions and Show Consequences to Raise Sustainability Awareness

### v1

@@ Abstract.
Unsustainable behaviors are challenging to prevent due to their long-term, often unclear consequences. Games offer a promising solution by creating artificial environments where players can immediately experience the outcomes of their actions. To explore this potential, we developed EcoEcho, a GenAI-powered game leveraging multimodal agents to raise sustainability awareness. These agents engage players in natural conversations, prompting them to take in-game actions that lead to visible environmental impacts. We evaluated EcoEcho using a mixed-methods approach with 23 participants. Results show a significant increase in intended sustainable behaviors post-game, although attitudes towards sustainability only slightly improved. This finding highlights the potential of multimodal agents and action-consequence mechanics to effectively motivate real-world behavioral changes such as raising environmental sustainability awareness.
1, 2, 3
@@
Figure 1. EcoEcho: An AI-Driven Role-Playing Game for Promoting Sustainable Behavior Awareness. The player is guided to take different actions through four distinct NPCs (Media Center - Lisa, Union Office - Bob, Union Building - Security, Government Building - Jonathan, respectively). The red text represents prompts or dialogues generated by the AI system to guide the player to engage in unsustainable behaviors. The blue text reflects the player’s expression of intent for clean energy within the game.
@@ 1. Introduction
Preventing unsustainable human behaviors has long been a challenge due to humanity’s inherent predisposition towards actions driven by short-term individual interests, coupled with the ambiguous nature of the long-term consequences associated with unsustainable practices  (Gual and Norgaard, 2010; Damasio, 1994) . Merely emphasizing the importance of sustainability fails to engender a visceral understanding among individuals. While humans are most profoundly influenced by experiencing the adverse outcomes or even failures resulting from their own actions, replicating such scenarios in the real world is neither feasible nor sufficiently protective of the environment. Moreover, individual differences in personality and attitudes significantly influence one’s receptiveness to sustainability goals, ultimately shaping their engagement with and commitment to sustainable behaviors.
@@
The advent of Generative AI (Feuerriegel et al., 2024), particularly the advanced natural language processing capabilities of language models, offers new avenues for simulating the harmful consequences of human behavior (Park et al., 2023; Link et al., 2016; Holmgard et al., 2014). We built EcoEcho, a virtual game environment that synthesizes visual, auditory, and narrative modalities to immerse players in the role of a conflicted scientist whose misinterpretation of his father’s dying wish leads him to impede clean energy advancement. To drive the plot forward, we harness GenAI’s ability to elicit natural language responses from players and seamlessly translate those verbal inputs into corresponding in-game actions. Through dynamic interactions with various stakeholders, such as journalists, union leaders, and ministers of energy, players engage in unsustainable behaviors within the game world, turning conversational choices into tangible consequences. By experiencing the negative outcomes of these actions, the game aims to increase awareness of sustainability issues and promote behavioral change both in-game and in real life.
@@
  • (c) Empirical Study: We assessed changes in players’ sustainability attitudes and behaviors before and after gameplay using surveys and semi-structured interviews. We found a significant increase in intended sustainable behaviors post-game, though attitudes only slightly improved.
@@ 2. Background
In section 2.1, we review the use of serious games for promoting sustainability awareness and education, given the increasing urgency of sustainability issues. We highlight how these games enhance engagement and knowledge, as well as the challenges in this field. In section 2.2, we review the role of generative AI to be a potential solution to address these challenges. We discuss how LLMs enhance game experiences by generating dynamic content and simulating realistic scenarios. Additionally, we also highlight some challenges to be noticed.
2.1. Games and Sustainability Awareness
@@
According to the Brundtland Report, sustainable development is defined as ”development that meets the needs of the present without jeopardizing the ability of future generations to meet their own needs” (Keeble, 1988). Building on this philosophy, the United Nations (UN) established the Sustainable Development Goals (SDGs), which prioritize essential areas such as energy utilization, water, and nourishment (Fund, 2015). The growing urgency of sustainability issues is gaining increasing attention (Owusu and Asumadu-Sarkodie, 2016; Harrington, 1992; Mihelcic et al., 2003), leading to the need for the development and integration of new educational applications designed to promote sustainable practices and sustainable awareness among the public (Altomonte et al., 2016). Games, especially serious games, are believed to serve as interactive tools to educate players on the complexities of sustainability, encouraging behavioral changes through engaging and immersive experiences (Stanitsas et al., 2019; Douglas and Brauer, 2021). Various types of games have been developed with distinct objectives and target audiences to facilitate sustainability transition. For instance, ”Keep Cool” (Eisenack, 2013) is a board game that simulates climate negotiations, helping players understand the impacts of different climate policies. Similarly, ”River Basin Game” (Lankford and Watson, 2007) is designed to teach water resource management by allowing players to manage and allocate water in a simulated river basin.”World Without Oil” (Rusnak et al., 2008) an alternate reality game, immerses players in a scenario where they must navigate a world facing an oil crisis, fostering awareness of energy dependency and sustainability. Additionally, digital games like ”About That Forest” (Games4Sustainability, nd) and ”Fate of the World” (Fate of the World Fandom, nd) task players with making decisions to mitigate climate change, balancing economic, social, and sustainable factors.
Research has shown that games could enhance motivation and engagement, leading to better knowledge outcomes in sustainability education. For instance, prior studies have found that games like ”Enercities,” (Knol and De Vries, 2011) which focuses on sustainable city planning, have been shown to increase awareness and understanding of sustainable development among players. ”EcoChains,” (Lee et al., 2016) a card game about Arctic marine food webs, helps players grasp the complexities of ecosystem interdependencies and the impact of climate change on biodiversity. ”The UVA Bay Game,” (Learmonth et al., 2011) a multiplayer simulation RPG focused on water management, has players assume roles such as policymakers, farmers, or developers, teaching them about watershed management through interactive decision-making and negotiation. These evaluations indicate that such games can be powerful educational tools, fostering not only knowledge acquisition but also motivation and engagement in sustainability topics.
@@
However, challenges remain in balancing engagement with educational value. Maintaining the educational purpose without compromising the fun and engaging aspects of the game is particularly challenging (Katsaliaki and Mustafee, 2015). Furthermore, developing high-quality serious games is resource-intensive, requiring time, expertise and funding, which limits the number and scope of games that can be produced (Stanitsas et al., 2019).
@@
In this study, we aimed to leverage the natural language understanding and generation capabilities of Large Language Models as the core gameplay mechanism in our game development. Additionally, we replaced traditional game programming with prompt engineering, using generative AI to produce the art and music assets within the game. Our goal is to explore a low-code, rapid development approach for creating AI-driven games on sustainability themes, providing a workflow for educators, non-profits, and other non-programmers to create serious games with purposeful objectives.
@@ 3.1. Game Design
Figure 2. The Game Overview (Exploring Unsustainable Decision Consequences): A Seven-Stage Game Framework from Opening to Final Decision, Including NPC Interaction, Time Travel, In-Game Assessments, and Multiple Endings(Bad Ending & Alternative Ending).
@@ 3.1.2. Game Narrative Design
The main story of the game, inspired by Back to the Future (Wikipedia contributors, 2024), subverts the traditional narrative found in sustainability games, where players typically perform good actions to achieve a sustainable world. In our game, EcoEcho, the story is set in the near future where sustainable energy is already widespread. Players control the character KI, who travels back to the past (our present world), and take actions to hinder the development of sustainable energy for personal reasons. If players simply follow the story without thinking deeply, they will ultimately reach a bad ending—where the Earth, overwhelmed by human pollution, becomes uninhabitable for humanity. This serves as a metaphor for the concept of the Anthropocene (Steffen et al., 2011). If we refuse to cultivate a sustainable energy mindset now, humanity may face a catastrophic future. Through the narrative, we connect players’ in-game experiences with their daily lives, encouraging them to transform their in-game emotions into a sustainable attitude in the real world, aligning with the concept of ”Reflection.”
@@
Figure 3. Visual Representations of Future Worlds: A) A Future World Where Clean Energy is Widespread B) A Future Where Traditional Energy Factories Still Exist C) A Future No Longer Suitable for Human Habitation
Figure 4. Overview of NPC Roles and Task Hierarchy: Protagonist Ki (a young scientist responsible for advancing the main storyline), Narrator Kane (an astronaut scientist and Ki’s father, providing story background and key emotional elements), Evaluator Emilia (a sustainability scientist and functional NPC, assessing player behavior and attitudes), Journalist Lisa (a reporter, advancing the story after completing tasks, Level 1), Security Guard (guiding players to the next storyline after completing specific tasks, Level 2), Union Leader Bob (driving key plot tasks, Level 3), and Minister of Energy Johnathan (the highest-level trigger NPC, final executor of key tasks, Level 4).
@@ 3.1.3. Gameplay Design
The core of the gameplay involves turning players’ words into actions, which then further impact the world. In traditional text-based adventure games, players typically make decisions by selecting from predefined options. However, in EcoEcho, we implement a natural language dialogue system, allowing players to engage in free-form conversations with Non-Player Characters (NPCs). By designing open conversations powered by language models, we aim to give players a strong sense of autonomy, enabling them to feel fully in control of their interactions and decisions. Autonomy, akin to the concept of ”Choice,” refers to the feeling of having control over one’s own decisions and actions. According to Self-Determination Theory, this can enhance motivation and engagement in games (Deci and Ryan, 2012; Ryan and Deci, 2000).
In the game, players must interact with NPCs to uncover their background stories and motivations. The player’s task is to skillfully craft their character’s dialogue to create prompts that drive the NPCs to take specific actions. For example, in level 1, the player must meet the demands of the senior journalist, Lisa, to progress. During their conversation, Lisa frequently emphasizes her desire for a big scoop, hinting that the player needs to share some future-world secrets about T-energy that the player already knows to gain her trust. Once this trust is established, the player is able to then influence Lisa’s attitude toward sustainable energy through dialogue to secure her support and assistance in halting the development of sustainable T-energy.
As the game progresses, inspired by the concept of ”Information,” we use environmental storytelling (Dubbelman, 2016) to subtly convey the impact of the player’s actions on the real world. Each time the player’s character returns to the future world, they can observe through the room’s window how the world environment deteriorates due to their actions (see Figure 3). This setup mirrors how the actions of human beings in the real world can actually influence the future. Overall, in this game, just as players use words to prompt NPCs’ actions, on a broader scale, the game’s overarching goal also serves as a prompt to drive the player’s own actions. However, unlike other language models that strictly follow player prompts, we encourage players to engage in critical thinking about AI prompts and make their own decisions at key moments. At the end of the game, we present a pivotal question asking the player whether they truly want to push for the repeal of the T-energy Development Act. If the player follows the game’s initially set objective and chooses ”yes,” they will reach the bad ending. However, if the player defies the game’s objective and chooses to abandon their previous efforts, they will achieve a good ending.
@@ 3.1.4. In-game Assessment Design
To assess players’ attitudes toward energy and sustainability within the game, we converted a 5-point Likert scale question on support for sustainable energy policies into an interactive scenario with a scientist character (see Figure 5). In this scenario, the scientist invites players to vote in a citizen petition, with each player receiving five votes. They can choose to cast between 1 and 5 votes. By having the scientist NPC interact with players at various stages of the game, we aim to track their attitudes toward energy and sustainability in the whole game process.
Figure 5. In-Game Assessment Scenarios: Appearing Four Times Post-NPC Dialogues with a 0-5 Voting Range, Example: ”In the first phase, we collected 587 signatures, nearing the goal of 618. How many will you contribute? (0-5)”
@@ 3.2. Gameful Human-Agent Interactions
Figure 6. Three-Stage Process of Prompt Engineering in Role-Playing Dialogue Systems: (1) Input Generation: Character Background, Details, Traits, and Motivation; (2) Prompt Summarization: Role, Objectives, Response Guidelines, and Instructions; (3) Response Generation: Llama-70B Output for Personalized Intelligent Dialogue.
@@ 3.2.1. NPC Prompt Engineering
Our prompt engineering involved employing prompt engineering with the Llama 3.1 70B model to simulate NPC roles. The overall process can be divided into three main phases (see Figure 6). We provide the complete prompt designs for NPCs in Appendix Character Prompts.

### v4

@@ Abstract.
Unsustainable behaviors are challenging to prevent due to their long-term, often unclear consequences. Games offer a promising solution by creating artificial environments where players can immediately experience the outcomes of their actions. To explore this potential, we developed EcoEcho, a GenAI-powered game leveraging multimodal agents to raise sustainability awareness. These agents engage players in natural conversations, prompting them to take in-game actions that lead to visible environmental impacts. We evaluated EcoEcho using a mixed-methods approach with 23 participants. Results show a significant increase in intended sustainable behaviors post-game, although attitudes towards sustainability only slightly improved. This finding highlights multimodal agents and action-consequence mechanics to effectively raising sustainability awareness and the potential of motivate real-world behavioral changes.
Multimodal Agents, Sustainability Awareness, Generative AI, In-Game Action
@@
Figure 1. EcoEcho: An AI-Driven Role-Playing Game for Promoting Sustainable Behavior Awareness. The player engage in a sequential interaction with four distinct Non-Player Characters (Media Center - Lisa, Union Office - Bob, Union Building - Security, Government Building - Jonathan, respectively) and guided to take different actions. The red text represents prompts or dialogues generated by the AI system to guide the player to engage in unsustainable behaviors. The blue text captures the player’s active expression of intent toward clean energy solutions.
@@ 1. Introduction
Preventing unsustainable human behaviors has long been a challenge due to humanity’s inherent predisposition towards actions driven by short-term individual interests, coupled with the ambiguous nature of the long-term consequences associated with unsustainable practices  (Gual and Norgaard, 2010; Damasio, 1994). Merely emphasizing the importance of sustainability fails to engender a visceral understanding among individuals. While humans are most profoundly influenced by experiencing the adverse outcomes or even failures resulting from their own actions, replicating such scenarios in the real world is neither feasible nor sufficiently protective of the environment. Moreover, individual personality differences significantly shape sustainability awareness—the recognition of the environment’s fragility and the need for its protection (Raymundo et al., 2019)—which drives engagement in and commitment to sustainable behaviors.
@@
The advent of Generative AI (Feuerriegel et al., 2024), particularly the advanced natural language processing capabilities of language models, offers new avenues for simulating the harmful consequences of human behavior (Park et al., 2023; Link et al., 2016; Holmgard et al., 2014). We developed EcoEcho, a virtual game environment designed to enhance players’ sustainability awareness. By integrating visual, auditory, and narrative modalities, EcoEcho immerses players in the role of a conflicted scientist whose misinterpretation of his father’s dying wish leads him to impede clean energy advancement. To drive the plot forward, we harness GenAI’s ability to elicit natural language responses from players and seamlessly translate those verbal inputs into corresponding in-game actions. Through dynamic interactions with various stakeholders, such as journalists, union leaders, and ministers of energy, players engage in unsustainable behaviors within the game world, turning conversational choices into tangible consequences. By experiencing the negative outcomes of these actions, the game aims to increase sustainability awareness both in-game and in real life.
@@
  • (c) Empirical Study: We assessed changes in players’ sustainability awareness through behaviors and attitudes before and after gameplay using surveys and semi-structured interviews. We found a significant increase in intended sustainable behaviors post-game, though sustainability attitudes only slightly improved.
@@ 2. Background
In section 2.1 and  2.2, we review the use of serious games for promoting sustainability awareness and education, given the increasing urgency of sustainability issues. We highlight how these games enhance engagement and knowledge, as well as the challenges in this field. In section 2.3, we review the role of generative AI to be a potential solution to address these challenges. We discuss how LLMs enhance game experiences by generating dynamic content and simulating realistic scenarios. Additionally, we also highlight some challenges to be noticed.
2.1. Serious Games in Sustainability: Theoretical Foundations
@@
According to the Brundtland Report, sustainable development is defined as ”development that meets the needs of the present without jeopardizing the ability of future generations to meet their own needs” (Keeble, 1988). Building on this philosophy, the United Nations (UN) established the Sustainable Development Goals (SDGs), which prioritize essential areas such as energy utilization, water, and nourishment (Fund, 2015). The growing urgency of sustainability issues is gaining increasing attention (Owusu and Asumadu-Sarkodie, 2016; Harrington, 1992; Mihelcic et al., 2003), leading to the need for the development and integration of new educational applications designed to promote sustainable practices and awareness among the public (Altomonte et al., 2016).
Games, particularly serious games, are believed to serve as interactive tools to educate players on the complexities of sustainability, encouraging behavioral changes through engaging and immersive experiences (Stanitsas et al., 2019; Douglas and Brauer, 2021). Douglas and Brauer (Douglas and Brauer, 2021) suggest that traditional serious games fit the SHIFT model (White et al., 2019), which encourages pro-environmental actions. The SHIFT framework integrates five psychological factors: social influence, habit formation, individual self, feelings and cognition, and tangibility. It leverages social norms and identity to drive behavior change, supports habit-building through incentives and feedback, emphasizes personal benefits of sustainability, balances emotional appeals with clear information, and makes abstract environmental impacts more tangible. According to the SHIFT model, serious games use interactive gameplay to allow players to practice sustainable actions, with the potential to turn these behaviors into long-term habits and positively influence sustainable attitudes.
2.2. Applications and Challenges of Serious Games in Sustainability
Various types of serious games have been developed with distinct objectives and target audiences to facilitate sustainability transition. For example, ”Keep Cool” (Eisenack, 2013) is a board game that simulates climate negotiations, helping players understand the impacts of different climate policies. Similarly, ”River Basin Game” (Lankford and Watson, 2007) is designed to teach water resource management by allowing players to manage and allocate water in a simulated river basin. ”World Without Oil” (Rusnak et al., 2008), an alternate reality game, immerses players in a scenario where they must navigate a world facing an oil crisis, fostering awareness of energy dependency and sustainability. Digital games such as ”About That Forest” (Games4Sustainability, nd) and ”Fate of the World” (Fate of the World Fandom, nd) task players with making decisions to mitigate climate change, balancing economic, social, and sustainable factors. Similarly, studies have shown that games like ”Enercities” (Knol and De Vries, 2011), focused on sustainable city planning, effectively increase players’ awareness and understanding of sustainable development. ”Climate Connected: Outbreak” (Fernández Galeote et al., 2023) is a VR game that situates its narrative in both fictional and real-world settings, offering a tailored experience that fosters learning, affective engagement, and motivation for real-world climate action. Another example, ”EcoChains” (Lee et al., 2016), a card game about Arctic marine food webs, helps players grasp the complexities of ecosystem interdependencies and the impact of climate change on biodiversity. Similarly, ”The UVA Bay Game” (Learmonth et al., 2011), a multiplayer simulation RPG, assigns players roles such as policymakers, farmers, or developers, teaching them about watershed management through interactive decision-making and negotiation.
@@ 2.2. Applications and Challenges of Serious Games in Sustainability
Despite these promising applications, challenges remain in balancing engagement with educational value. Maintaining the educational purpose without compromising the fun and engaging aspects of the game is particularly challenging (Katsaliaki and Mustafee, 2015). Some serious games feature overly complicated mechanisms, resulting in steep learning curves that primarily appeal to individuals already interested in the domain or key stakeholders, thereby failing to engage a broader audience. Additionally, certain serious games rely heavily on theoretical education rather than interactive experiences, which can hinder their ability to sustain learners’ interest. Moreover, developing high-quality serious games is resource-intensive, requiring time, expertise, and funding, which limits the number and scope of games that can be produced (Stanitsas et al., 2019).
@@
In this study, we explored the use of Large Language Models (LLMs) for natural language understanding and generation as the core gameplay mechanism in our game. Compared to traditional serious games with complex mechanics that require significant learning effort, our approach allows players to interact with NPCs only simply using natural language, reducing the learning curve and enabling even non-gamers to quickly get started. Additionally, we replaced traditional game programming with prompt engineering, using generative AI to produce the art and music assets within the game. Our goal is to explore a low-code, rapid development approach for creating AI-driven games on sustainability themes, providing a workflow for educators, non-profits, and other non-programmers to create serious games with purposeful objectives.
@@ 3. System Design
Figure 2. Sequential visualization of the game’s narrative structure across seven key stages: (1) game initiation, (2) opening sequence introducing the protagonist’s world, (3) NPC interactions, (4) time travel mechanics, (5) player assessment phases, (6) critical decision point, and (7) branching endings (Alternative or Bad) based on player choices regarding environmental sustainability.
@@ 3.1. Game Design
Figure 3. Overview of NPC Roles and Task Sequence Hierarchy: Protagonist Ki (a young scientist responsible for advancing the main storyline), Narrator Kane (an astronaut scientist and Ki’s father, providing story background and key emotional elements), Evaluator Emilia (a sustainability scientist and functional NPC, assessing player behavior and attitudes), Journalist Lisa (a reporter, advancing the story after completing tasks, Level 1), Security Guard (guiding players to the next storyline after completing specific tasks, Level 2), Union Leader Bob (driving key plot tasks, Level 3), and Minister of Energy Johnathan (the highest-level trigger NPC, final executor of key tasks, Level 4).
@@ 3.1.2. Game Narrative Design
The main story of the game, inspired by Back to the Future (Wikipedia contributors, 2024), subverts the traditional narrative found in sustainability games, where players typically perform good actions to achieve a sustainable world or save a polluted world (Fernández Galeote et al., 2023). In our game, EcoEcho, the story is set in the near future where sustainable energy is already widespread. Players control the character KI, who travels back to the past (our present world), and take actions to hinder the development of sustainable energy for personal reasons. If players simply follow the story without thinking deeply, they will ultimately reach a bad ending—where the Earth, overwhelmed by human pollution, becomes uninhabitable for humanity. This serves as a metaphor for the concept of the Anthropocene (Steffen et al., 2011). If we refuse to cultivate a sustainable energy mindset now, humanity may face a catastrophic future. Through the narrative, we connect players’ in-game experiences with their daily lives, encouraging them to transform their in-game emotions into a sustainable attitude in the real world, aligning with the concept of ”Reflection.”
@@ 3.1.3. Gameplay Design
Main Game Mechanism: The core of the gameplay involves turning players’ words into actions, which then impact the world. In traditional text-based adventure games, players make decisions by selecting predefined options. However, in EcoEcho, a natural language dialogue system enables free-form conversations with Non-Player Characters (NPCs). This design empowers players with a strong sense of autonomy, allowing them to feel fully in control of their interactions and decisions. Autonomy, akin to the concept of ”Choice,” refers to the feeling of having control over one’s own decisions and actions. According to Self-Determination Theory, this can enhance motivation and engagement in games (Deci and Ryan, 2012; Ryan and Deci, 2000).
Player-NPC Interaction: Players must interact with NPCs to uncover their background stories and motivations. The gameplay revolves around crafting the player character’s dialogue to create prompts that drive NPCs to take specific actions. For example, in level 1, the player meets a senior journalist, Lisa, whose support is crucial for progression. During their conversation, Lisa frequently emphasizes her desire for a big scoop, hinting that the player needs to share future-world secrets about T-energy to gain her trust. Once trust is established, the player can influence Lisa’s attitude toward sustainable energy, securing her assistance in halting the development of sustainable T-energy.
Figure 4. Parallel views of three potential futures: (A) a world transformed by clean energy adoption, (B) a landscape dominated by traditional energy infrastructure, and (C) an environment rendered uninhabitable by environmental degradation.
Environmental Impact as Feedback: Inspired by the concept of ”Information,” EcoEcho employs environmental storytelling (Dubbelman, 2016) to convey the consequences of the player’s actions on the real world. Each time the player character returns to the future world, they can observe the environmental deterioration through the room’s window. This mirrors how human actions in the real world influence the future. The evolving environment serves as visual feedback, linking the player’s decisions to tangible outcomes in the game world.
Critical Thinking and Endgame Decision: The overarching goal of EcoEcho encourages players to reflect critically on AI prompts and their own decisions. Unlike other language-based games that strictly follow player prompts, EcoEcho challenges players to think independently at key moments. At the game’s conclusion, players face a pivotal decision: whether to push for the repeal of the T-energy Development Act. If the player follows the game’s initial objective and chooses ”yes,” they will reach the bad ending. Conversely, defying the game’s objective by abandoning previous efforts leads to the good ending. This twist reinforces the theme of critical decision-making and its broader implications.
@@ 3.1.4. In-game Assessment Design
To assess players’ attitudes toward energy and sustainability within the game, we converted a 5-point Likert scale question on support for sustainable energy policies into an interactive scenario with a scientist character (see Figure LABEL:fig:fig4). In this scenario, the scientist invites players to vote in a citizen petition, with each player receiving five votes. They can choose to cast between 1 and 5 votes. By having the scientist NPC interact with players at various stages of the game, we aim to track their attitudes toward energy and sustainability in the whole game process.
@@ 3.2. Gameful Human-Agent Interactions
Figure 5. Architectural overview of a three-stage prompt engineering framework for role-playing dialogues: initial character profile compilation, prompt synthesis and structuring, and large language model (Llama-70B) dialogue generation, illustrated through an investigative journalist character scenario.
@@ 3.2.1. NPC Prompt Engineering
Our prompt engineering involved employing prompt engineering with the Llama 3.1 70B model to simulate NPC roles. The overall process can be divided into three main phases (see Figure 5). We provide the complete prompt designs for NPCs in Appendix B.